OpenStudy (anonymous):

If today is thursday, what is 550 days from today? Is it monday?

OpenStudy (turingtest):

550/7=?

OpenStudy (anonymous):

a biiiig number

OpenStudy (anonymous):

78.6

OpenStudy (anonymous):

round it up.

OpenStudy (anonymous):

up to what?

OpenStudy (anonymous):

79

OpenStudy (turingtest):

leave it in mixed fraction form. That is, leave the remainder visible.

OpenStudy (turingtest):

for instance 29/5=5+4/5 so do that with 550/7...

OpenStudy (turingtest):

just do long division. what is the remainder?

OpenStudy (anonymous):

For such a problem, you need to do a mod. So, 550 mod 7. The value gives you the days since Thursday, which should give you the exact day.

OpenStudy (anonymous):

so mu cuzzin said dont ask him how he know but its a monday he is a smart a**

OpenStudy (anonymous):

550 mod 7 = 4. So, it is Monday.

OpenStudy (anonymous):

June 27, 2013

OpenStudy (anonymous):

thursday

OpenStudy (anonymous):

It's very easy...simply take some calculations. If today is december 26 monday on my side then 550 days after it will be 28 june 2013 (Friday). Similarly if today is december 29 thursday 2011 then after 550 days it will be 1st JULY 2013 (Monday).

OpenStudy (anonymous):

So yes It is monday! :)

OpenStudy (anonymous):

Theta45 - use the simple math of a modulus operation. You are complicating it. :)

OpenStudy (anonymous):

you can divide 550 by 7, it gives you 78.6, so you multiply 7*78 and it gives you 546, then you add four days so you get to 550, and four days after thursday is monday. so monday is the answer